OpenStudy (anonymous):

-2x^2+8x+3=1 <--- solve by the quadratic formula , set the equation to "0" first. must be in simplest radical form, show work please

OpenStudy (anonymous):

-2x^2+8x+3=1 -2x^2 + 8x +2 = 0 x = 2-sqrt(6) x = 2+sqrt(6)